OpenStudy (anonymous):

plz help simplify the given expression sqrt-108+sqrt-27

OpenStudy (anonymous):

start with \[\sqrt{36\times 3\times (-1)}+\sqrt{9\times 3\times (-1)}\] then go to \[\sqrt{36}\times \sqrt{3}i +\sqrt{9}\sqrt{3}i\] then \[6\sqrt{3}i+3\sqrt{3}i\] and finally add them up

OpenStudy (anonymous):

9sqrt3i is my answer

OpenStudy (anonymous):

that would be right

OpenStudy (anonymous):

ok than you for helping me